If x =-2, y=-4/5 and z =-1/2 show that x,y,z are 3 consecutive of a harmonic progression
Answer by richwmiller(17219)

If you mean x =-2, y=-4/5 and z =-1/2
then
-1/2,-5/4 , -2
-.5 ,-1.25, -2
add -.75 or 3/4
so the reciprocals are in a AP with the d being 3/4